OpenStudy (anonymous):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 60 dumpster loads of waste removed. In week 10, there were 40 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. f(x) = - 4x + 60 f(x) = 4x + 60 f(x) = - 4x + 80 f(x) = 4x + 80

OpenStudy (anonymous):

i know its either A or B but i dont know how they got 60 or 80 please help me!

OpenStudy (mitu12):

i was thinking it would be c

OpenStudy (anonymous):

i meant C not B XD

OpenStudy (anonymous):

because the 4 is negative so that eliminates B and D but now what do i do?

OpenStudy (mitu12):

y = -4x + b 60 = -4(5) + b solve b

OpenStudy (anonymous):

thats what i cant figure out!

OpenStudy (anonymous):

i dont know how to find b tho i have y=-4x

OpenStudy (anonymous):

just missing that b but idk what to do next :l

OpenStudy (mitu12):

60=−4(5)+b 60=−20+b 60−(−20)=b 80=b

OpenStudy (anonymous):

YOU'RE RIGHT!

OpenStudy (mitu12):

therefore the equation is c

OpenStudy (anonymous):

thanks alot man and also thanks for actually showing it not just giving the answer! need to know this stuff :3

OpenStudy (mitu12):

no problem glad i could help!